219 199 232 183 217 222 205 199 239 233 241 205 214 198 250 245 211 200 259 222 239 235 199 247 228 242 228 238 227 204 211 300 280 288 271 111 132 156 190 198 245 232 241 243 280 265 276 290 227


group the data and construct a contigency table containing class limit,class boundaries,frequency,midpoint,relative frequency,less than cumulative frequency and more than cumulative frequency.use the interval of 111-138, 139-166
1
Expert's answer
2020-09-21T16:07:10-0400


Less than C.F More than C.F

Class "f" Class Midpoint Relative Marks C.F Marks C.F

Limit Boundaries Frequency

111-138 2 110.5-138.5 124.5 2/49 < 138.5 2 "\geq" 110.5 49

139-166 1 138.5-166.5 152.5 1/49 < 166.5 3 "\geq" 138.5 47

167-194 2 166.5-194.5 180.5 2/49 < 194.5 5 "\geq" 166.5 46

195-222 16 194.5-222.5 208.5 16/49 < 222.5 21 "\geq" 194.5 44

223-250 19 222.5-250.5 236.5 19/49 < 250.5 40 "\geq" 222.5 28

251-278 4 250.5-278.5 264.5 4/49 < 278.5 44 "\geq" 250.5 9

279-306 5 278.5-306.5 292.5 5/49 < 306.5 49 "\geq" 278.5 5

49